Namclear agency is a body tasked with the clearance of all bank guaranteed cheques from all commecial banks in the country. According to the agency it was observed that 1 in 20 cheques in Namibia are fraudulent. suppose you radomly selected 20 cheques from this population of cheques.
1 what is the probability that no cheque in the samle will be fraudulent?
2 what is the probability that one cheque will in the sample will fraudulent?
3 what is the probability that one or fewer cheques in the sample will be fraudulent?
4 what is the probability that at least 2 cheques in the sample will be fraudulent?

Answer by stanbon(75887)

Namclear agency is a body tasked with the clearance of all bank guaranteed cheques from all commecial banks in the country. According to the agency it was observed that 1 in 20 cheques in Namibia are fraudulent. suppose you radomly selected 20 cheques from this population of cheques.
P(bad) = 1/20 ; P(good) = 19/20
------------------------------------------
1) what is the probability that no cheque in the sample will be fraudulent?
P(all good) = (19/20)^20 = 0.3585
--------------------------------------------
2 what is the probability that one cheque will in the sample will fraudulent?
P(x = 1)= 20C1*20)(19/20)^19 = 20*(1/20)*(19/20)^19 = 0.00000000908....
---------------------
3 what is the probability that one or fewer cheques in the sample will be fraudulent?
P(x = 1 or 0) = P(1) + P(0) = 0.35850000908
---------------------------------------------------
4 what is the probability that at least 2 cheques in the sample will be fraudulent?
P(2 <= x <= 20) = 1 - P(x = 1 or x= 0) = 1- 0.35950000908 - 0.64150...
========================
Cheers,
Stan H.